# Gravitational wave probes of particle dark matter: a review

### v3

@@ III.4.2 Variation of fundamental constants constraints
The strongest constraints from pulsar timing arrays are on the couplings to the electron and muons. Experiments that directly use atomic clocks to search for ultralight dark matter are insensitive to the electron coupling, since changing the electron mass does not affect the spacing between energy levels Kaplan et al. (2022). In terms of (the lack of) other constraints on muons, laboratory experiments do not have enough muons to harness on earth to study the coupling of DM to them, while, in contrast, neutron stars have an abundance of muons.
@@
In the gray-shaded region, the interaction between gravity and DM is stronger, and thus constraints on ultralight dark matter coupling to standard-model particles alone would not be valid in that region. Furthermore, these constraints assume that scalar particles constitute all of DM, which may not be true. If they only comprise a fraction $f$ of DM, then the constraints will be weakened by $\sqrt{f}$, because of the $\sqrt{\rho_{\text{DM}}}$ term in the amplitude of these interaction signals.
Figure 10: Taken from Afzal et al. (2023). Upper limits from an analysis of NANOGrav data on the coupling of ultralight dark matter to various standard-model particles: quarks, electrons, muons, photons and gluons, shown in red for the correlated case, and red dashed for the uncorrelated case. In this model, DM would alter the moment of inertia of neutron stars and thus affect the times of arrivals of pulses in pulsar timing arrays. The constraints on one coupling constant assume that all other coupling constants are set to 0. The black dashed line, and the gray shaded region, indicate where the amplitude of the signal would be less than the amplitude of the signal arising from DM coupling only to gravity. Current constraints “Rb/Cs atomic clocks” (purple) are from Hees et al. (2016), “Al/Hg atomic clocks” (turquoise) are from Beloy et al. (2021), “MICROSCOPE” (teal) are from Bergé et al. (2018), “H/Si clock shift” (orange) are from Beloy et al. (2021), and “NS binary system” are from Kumar Poddar et al. (2019) and Dror et al. (2020).
The scalings of these constraints with the DM mass should be discussed. At high masses, the signal is essentially deterministic, and the signal-to-noise ratio scales as the square of the signal amplitude, which means that $d_{i}\propto m^{2}$. At low masses ($m<1/T_{\text{obs}})$, the signal is observed in different coherence patches, since the coherence length is shorter than the pulsar separation. The signal is no longer a sinusoid but can be expressed as a polynomial in $mt$. The first two terms of this polynomial ($mt$ and $(mt)^{2}$) are degenerated with pulsar timing terms, and thus the first observable term is $(mt)^{3}$, meaning that the signal-to-noise ratio is proportional to $m^{3}$. Thus, $d_{i}\propto 1/m$, which explains the scaling at low masses. Still, though, red noise causes the constraints to flatten as the mass decreases, which, at first, hides this scaling.
III.4.3 Constraints on vector dark matter
@@
Fig. 11 shows constraints on the coupling constants $g_{B}=\epsilon_{B}$ and $g_{B-L}=\epsilon_{B-L}$ as a function of the DM mass and oscillation frequency. Again, the gray-shaded region indicates the regime in which the coupling of DM to gravity becomes stronger than the coupling of DM to standard-model particles. We can see that the NANOGrav constraints greatly surpass those that come from MICROSCOPE at very low masses. Furthermore, these constraints assume that vector bosons constitute all of DM, which may not be true. If they only comprise a fraction $f$ of DM, then the constraints will be weakened by $\sqrt{f}$, because of the $\sqrt{\rho_{\text{DM}}}$ term in the amplitude of these interaction signals.
@@
Figure 11: Taken from Afzal et al. (2023). Upper limits from analyses of NANOGrav pulsar timing array data on the coupling of DM to baryon and baryon-lepton number for the correlated (solid lines) an uncorrelated (dashed line) cases. In this case, dark photon DM would couple to standard model particles in the earth and the pulsars themselves. The black dashed line, and the gray shaded region, indicates where the amplitude of the signal would be less than the amplitude of the signal arising from DM coupling only to gravity. The constraints from tests of the equivalence principle (“MICROSCOPE”) are shown in teal (Bergé et al., 2018), and the constraints previously set by the PPTA Collaboration are reported in yellow (Xue et al., 2022).
The recent NANOGRAV evidence for a stochastic GW background is also consistent with a vector ultralight dark matter hypothesis. Ref. Chowdhury et al. (2024) has found that ultralight dark matter that couples to muons ($L_{\mu}-L_{\tau})$ is actually more consistent with NANOGRAV’s observations than a stochastic GW background. This type of ultralight dark matter would cause the pulsars to oscillate (because they are composed of muons) and induce a spin-dependent frequency shift in the pulses that would change the angular correlations with respect to the stochastic GW background case (because of the gravitational potential of the ultralight dark matter). The mass and coupling constants of this ultralight dark matter are not ruled out by any other experiment, and lie in the region
III.4.4 Constraints on conformal DM
DM that conformally (universally) couples to gravity could alter the moment of inertia of rotating pulsars via a new gravity-mediated force between neutron stars and ultralight dark matter Kuntz and Barausse (2024), which is called the Nordtvedt effect Nordtvedt (1968). If such a coupling exists, the times of arrivals of pulsars would change in a deterministic way, and would also result in a violation of the equivalence principle. This model of conformal DM has been extensively constrained by looking for impacts on the timing solutions of measured binary pulsar Damour and Taylor (1992); Kramer et al. (2006).
Ref. Smarra et al. (2024) uses European pulsar timing array data to constrain linear (Fierz-Jordan-Brans-Dicke (FJBD) theory) Fierz (1956); Jordan (1959); Brans and Dicke (1961); Dicke (1962) and quadratic (Damour-Esposito-Farese (DEF) theory) Damour and Esposito-Farese (1992, 1993) conformal DM couplings to gravity by deriving how the pulsar times of arrivals would change in the presence of the ultralight dark matter fields. In the case of FJBD conformal coupling, the residuals would appear as:
$\Delta t(t)$ $=-\int\mathrm{d}t\frac{\delta\Omega_{\text{obs}}}{\bar{\Omega}}$
$=2\alpha s_{I}\left.\frac{\sqrt{\rho_{\text{DM}}}}{M_{\text{Pl}}m^{2}}\hat{\phi}(\bm{x})\sin(m_{\rm DM}t+\theta(\bm{x}))\right|_{t_{\text{start}}-\frac{d}{c}}^{t_{\text{end }}-\frac{d}{c}},$ (88)
where $s_{I}$ is the angular moment sensitivity parameter computed in Kuntz and Barausse (2024) and is of $\mathcal{O}(1)$ in FJBD theory, $\alpha$ is the strength of the coupling, and $\hat{\phi}$ is a stochastic parameter whose value depends on whether the correlated, uncorrelated or pulsar correlated scenarios are considered. $\theta(\mathbf{x})$ is a random phase.
In Fig. 12, we show the upper limits on the linear coupling parameter $\alpha$ using European pulsar timing array data. However, Ref. Smarra et al. (2024) notes that there is some evidence for additional signal power at two masses: $m\sim 10^{-22.7}$ eV and $m\sim 10^{-21.4}$ eV, but note that other physical processes could responsible for these excesses.
Figure 12: Taken from Smarra et al. (2023). Constraints on the coupling $\log_{10}\alpha$ of conformal DM to gravity at the 95% confidence-level (solid line) for the correlated case compared to the analytical estimate obtained by using the upper limits from the European pulsar timing array search in Smarra et al. (2023) to estimate $\alpha$ (brown dashed lines). $\rho_{\text{DM}}=0.4\,\text{GeV/cm}^{3}$. Here, this model of DM would change the moment of inertia of neutron stars in the used in pulsar timing arrays, resulting in changes of the times of arrival of the pulses. If the true density of DM that conformal DM composes is less than $\rho_{\text{DM}}$, the limits will suffer, which is shown by the dotted line, for values of $\rho$ given in Smarra et al. (2023). AP4 is a particular equation-of-state of neutron stars Akmal et al. (1998).
III.4.5 Constraints on axions through polarization measurements
Changes in polarization of light by axions could also be probed in pulsar timing array experiments Xue et al. (2024); Porayko et al. (2024). Just as in ground-based interferometers, axions could induce cosmic birefrigence, an oscillation of the polarization plane of a linearly polarized source. Since pulsars constantly emit light that is beamed at earth, its polarization can be measured over time, and any changes in it could be attributed to axion-like DM/photon interactions. In Fig. 13, we show upper limits on this coupling from the Parkes Pulsar Polarization Array (PPPA), the first search of its kind. The constraints here outperform existing experiments at low axion masses.
@@ III.4.5 Constraints on axions through polarization measurements
Figure 13: Taken from Xue et al. (2024). Upper limits from analyses of European pulsar timing array data on the coupling of axion-like DM to photons.$\rho_{\text{DM}}=0.4\,\textrm{GeV}/\textrm{cm}^{3}$. The blue and red curves correspond to the limits derived using the auto-correlation-only and full-correlation signal models, respectively. Excluded portions of the axion mass /coupling parameter space lie above these curves, and constraints from different experiments are also shown. Key timescales are shown with vertical dotted lines, included $T_{\text{obs}}=18$ years. The lower panel shows the Bayes factors for each of the scenarios.
III.5 Prospects for detecting the coupling of ultralight dark matter to the interferometers
@@
Though no searches so far have yielded conclusive evidence of any DM/standard model coupling with the interferometers, impressive constraints, relative to existing DM experiments, have been set on these interaction models. The current searches and their constraints have been framed only in terms of upper limits on dilatons, dark photons and DM that couples to gravity; however, they are actually sensitive to any kind of DM interaction that would cause a differential strain in the interferometers. In fact, to derive these constraints, we first compute the minimum detectable signal amplitude for a generic quasi-monochromatic signal as a function of frequency, and then map these values to constraints on the DM/standard model coupling constants. It is thus simply a matter of a few relations that these results are interpreted in terms of two relatively well-motivated DM models.
@@
Of course, ultralight dark matter could take on any mass a priori, and ground-based GW detectors only allow us to probe a small region of this parameter space, from $\sim 10^{-14}-10^{-12}$ eV. Space-based GW detectors will permit being sensitive to masses a few orders of magnitude smaller, while high-frequency gravitational-wave (HFGW) detectors could see DM particles a few orders of magnitude larger Aggarwal et al. (2021). However, it should be noted that when placing constraints on ultralight dark matter, the stochasisity of the signal could affect the limits: if we happen to “get unlucky” and observe the signal when the phase offset reduces the signal amplitude to zero, as can be seen at the zero points in Fig. 2, we would have no constraint. Such considerations were made in Nakatsuka et al. (2023); Boddy et al. (2025) for ground-based interferometers and pulsar timing arrays.
@@
The prospects for future ground-based GW interferometers, such as ET and Cosmic Explorer, to observe ultralight dark matter interactions are bright. Within a decade or so of observation, such detectors, currently a couple of orders of magnitude behind fifth-force constraints, could actually be competitive in this portion of the parameter space, thus permitting an extremely strong probe of new physics Pierce et al. (2018); Morisaki et al. (2021).
@@
Though GW interferometers can probe extremely small DM/standard model couplings, we do not have an estimate of a minimum strength of these interactions. We thus have no guarantee, in any search that we do, that we are getting closer to probing the true coupling value between ordinary matter and the dark sector. Such a caveat is balanced by the relatively model agnostic nature of our searches, which make minimal assumptions about the signal model, and essentially just look for correlated noise. While we probably cannot believe just one model for DM/standard model interactions, we can be sure that our techniques are sensitive to a wide range of theories, even ones that have yet to be thought of.
IV Macroscopic dark matter transiting through the interferometers
IV.1 Background
As mentioned in Section I, the mass of DM can span many orders of magnitude, including being “macroscopic”, that is, from being $M_{\rm DM}\sim[1,10^{9}]$ kg Hall et al. (2018). If DM exists in this mass range, it could transit through GW interferometers, and thus be searched for in LIGO, Virgo and KAGRA data. The event rate, however, depends on the mass of DM, and there are orders of magnitude fewer heavy DM particles in any given region of space than the number of ultralight dark-matter particles Hall et al. (2018):
$\rho_{\text{DM}}=M_{\text{DM}}n_{\text{DM}}~{}~{}\Longrightarrow~{}~{}\frac{L_{\rm DM}}{10^{4}~{}\text{km}}\simeq 1.2\times\left(\frac{M_{\text{DM}}}{1~{}\text{kg}}\right)^{1/3},$ (89)
where $n_{\rm DM}$ is the number density of DM, and $L_{\rm DM}$ is a characteristic length scale for the transiting DM. For 1-kg DM, the flux of DM through a given region of space is given by $\Phi\sim n_{\rm DM}v_{0}\sim 3\times 10^{-10}$ km-2 s-1, meaning that approximately one transient event could be observed per year if the impact parameter of such an event is 10 km, which is similar in size to the three or four kilometer-long arms of LIGO, Virgo and KAGRA.
IV.2 Yukawa-like interactions

### v4

@@ III.3.3 Variation of fundamental constants constraints
The strongest constraints from pulsar timing arrays are on the couplings to the electron and muons. This is in contrast to constraints that come from atomic clock experiments, which also aim to find ultralight dark matter but are insensitive to the electron coupling, since changing the electron mass does not affect the spacing between energy levels Kaplan et al. (2022). In terms of (the lack of) other constraints on muons, laboratory experiments do not have enough muons to harness on earth to study the coupling of DM to them, while, in contrast, neutron stars have an abundance of muons.
@@
In the gray-shaded region, the interaction between gravity and DM is stronger, and thus constraints on ultralight dark matter coupling to standard-model particles alone would not be valid in that region. Furthermore, these constraints assume that scalar particles constitute all of DM, but if they only comprise a fraction $f$ of DM, then the constraints will be weakened by $\sqrt{f}$, because of the $\sqrt{\rho_{\text{DM}}}$ term in the amplitude of these interaction signals.
Figure 10: Taken from Afzal et al. (2023). Upper limits from an analysis of NANOGrav data on the coupling of ultralight dark matter to various standard-model particles: quarks, electrons, muons, photons and gluons, shown in red for the correlated case, and red dashed for the uncorrelated case. In this model, DM would alter the moment of inertia of neutron stars and thus affect the ToAs of pulses in pulsar timing arrays. The physics behind this form of DM has been discussed in Section II.3.2. The constraints on one coupling constant assume that all other coupling constants are set to 0. The black dashed line, and the gray shaded region, indicate where the amplitude of the signal would be less than the amplitude of the signal arising from DM coupling only to gravity. Current constraints “Rb/Cs atomic clocks” (purple) are from Hees et al. (2016), “Al/Hg atomic clocks” (turquoise) are from Beloy et al. (2021), “MICROSCOPE” (teal) are from Bergé et al. (2018), “H/Si clock shift” (orange) are from Beloy et al. (2021), and “NS binary system” are from Kumar Poddar et al. (2019) and Dror et al. (2020). The shape of these constraints is similar to those in Fig. 11, since both searches look for nearly sinusoidal signals embedded in pulsar timing array data with similar noise characteristics, even though the underlying DM coupling mechanisms differ.
The scalings of these constraints with the DM mass should be discussed. At high masses, the signal is essentially deterministic, and the SNR scales as the square of the signal amplitude, which means that $d_{i}\propto m_{\rm DM}^{2}$. At low masses ($m_{\rm DM}<1/T_{\text{obs}})$, the signal is observed in different coherence patches, since the coherence length is shorter than the pulsar separation. The signal is no longer a sinusoid but can be expressed as a polynomial in $mt$, where $t$ is time. The first two terms of this polynomial ($m_{\rm DM}t$ and $(m_{\rm DM}t)^{2}$) are degenerated with pulsar timing terms, and thus the first observable term is $(m_{\rm DM}t)^{3}$, meaning that the SNR is proportional to $m_{\rm DM}^{3}$. Thus, $d_{i}\propto 1/m_{\rm DM}$, which explains the scaling at low masses. Still, though, red noise causes the constraints to flatten as the mass decreases, which, at first, hides this scaling.
III.3.4 Constraints on vector dark matter
@@
Fig. 11 shows constraints on the coupling constants $g_{B}=\epsilon_{B}$ and $g_{B-L}=\epsilon_{B-L}$ as a function of the DM mass and oscillation frequency. Again, the gray-shaded region indicates the regime in which the coupling of DM to gravity becomes stronger than the coupling of DM to standard-model particles. We can see that the NANOGrav constraints greatly surpass those that come from MICROSCOPE at very low masses. Furthermore, these constraints assume that vector bosons constitute all of DM, but if they only comprise a fraction $f$ of DM, then the constraints will be weakened by $\sqrt{f}$.
@@
Figure 11: Taken from Afzal et al. (2023). Upper limits from analyses of NANOGrav pulsar timing array data on the coupling of DM to baryon and baryon-lepton number, shown for the correlated (solid lines) and uncorrelated (dashed lines) cases. In this scenario, dark photon DM couples to standard model particles in both the Earth and the pulsars. The physics behind this form of DM has been discussed in Section II.4. The black dashed line and gray shaded region indicate where the amplitude of the signal would be less than that from gravitational coupling alone. Constraints from equivalence principle tests (“MICROSCOPE”) are shown in teal (Bergé et al., 2018), and previous limits from the PPTA Collaboration are shown in yellow (Xue et al., 2022). The similarity in shape between the curves in this figure and those in Fig. 10 arises because both searches target nearly sinusoidal signals embedded in pulsar timing array data with similar noise characteristics. Although the underlying DM models differ—scalar versus vector, moment of inertia changes versus direct forces—the observable effect in each case is the same: modulations in pulse ToAs. The coupling strength to standard-model particles determines the amplitude of these signals.
Furthermore, recent evidence for a stochastic GW background reported by NANOGrav has also been interpreted in terms of vector ultralight dark matter. In particular, Ref. Chowdhury et al. (2024) showed that a vector ultralight dark matter model coupled to muons ($L_{\mu}-L_{\tau}$) can provide an even better fit to NANOGrav’s observations than a conventional stochastic GW background. In this scenario, pulsars undergo oscillations due to their muon content, leading to a spin-dependent frequency shift in the pulses. This produces angular correlations that differ from the standard Hellings–Downs curve expected for a stochastic GW background, thereby offering a way to test the vector ultralight dark-matter hypothesis against future pulsar timing array data.
III.3.5 Constraints on axions through polarization measurements
As described in Section II.2, axion-like dark matter induces an oscillatory modulation of the polarization plane of linearly polarized light—cosmic birefringence—via its coupling to photons.
Recent analyses of European pulsar timing array data, particularly from the Parkes Pulsar Polarization Array (PPPA), have performed the first dedicated searches for these axion-induced polarization oscillations Xue et al. (2024); Porayko et al. (2025). By tracking polarization angles over long time baselines, these searches place upper limits on the axion-photon coupling constant, $g_{a\gamma}$, for axion masses in the ultra-light regime.
Fig. 12 shows the resulting constraints from this analysis. The red and blue curves correspond to different statistical treatments of the data, including models that account for spatial correlations between pulsars. Notably, these PTA constraints improve upon previous laboratory and astrophysical bounds at very low axion masses (below approximately $10^{-22}$ eV), where the long coherence time of the axion field aligns well with the decade-scale observation windows of PTAs.
These results highlight the complementary sensitivity of PTAs compared to ground-based interferometers, which probe higher axion masses but with comparable constraints on $g_{a\gamma}$. Together, they cover a broad mass range, making multi-messenger searches a powerful strategy for testing axion dark matter scenarios.
@@ III.3.5 Constraints on axions through polarization measurements
Figure 12: Taken from Xue et al. (2024). Upper limits from analyses of European pulsar timing array data on the coupling of axion-like DM to photons. $\rho_{\text{DM}}=0.4\,\textrm{GeV}/\textrm{cm}^{3}$. The physics behind this form of DM has been discussed in Section II.2. The blue and red curves correspond to the limits derived using the auto-correlation-only and full-correlation signal models, respectively. Excluded portions of the axion mass /coupling parameter space lie above these curves, and constraints from different experiments are also shown. Key timescales are shown with vertical dotted lines, included $T_{\text{obs}}=18$ years. Smoothed constraints are shown as dashed lines in cyan, yellow, green, and purple, representing limits from polarization angle (PA) variations observed in the Crab Nebula by POLARBEAR Adachi et al. (2024), analyses of PPTA DR1 and Crab Nebula data from QUIJOTE Castillo et al. (2022), measurements of CMB polarization by SPT-3G Ferguson et al. (2022), and studies of axion-like DM-induced washout effects during recombination from the Planck mission Fedderke et al. (2019), respectively. Additional constraints from the CAST experiment Anastassopoulos et al. (2017), as well as observations of SN1987A Payez et al. (2015) and X-ray spectral distortions in quasar H1821+643 Reynés et al. (2021), are shown as black dashed lines. The lower panel shows the Bayes factors (BFs, essentially, the preference for DM to be in the data versus noise) for each of the scenarios. The BF stays around zero, although some significant spikes are apparent, which has been found to be due to imprecise noise modelling, not astrophysical signals.
III.4 Prospects for detecting the coupling of ultralight DM to GW interferometers
@@
Though no searches so far have yielded conclusive evidence of any DM/standard model coupling with the interferometers, impressive constraints, relative to existing DM experiments, have been set on these interaction models. The current searches and their constraints have been framed only in terms of upper limits on dilatons, dark photons and DM that couples to gravity; however, they are actually sensitive to any kind of DM interaction that would cause a differential strain in the interferometers. In fact, to derive these constraints, we first compute the minimum detectable signal amplitude for a generic quasi-monochromatic signal as a function of frequency, as outlined in Section III.1, and then map these values to constraints on the DM/standard model coupling constants. It is thus simply a choice that these results are interpreted in terms of two relatively well-motivated DM models.
@@
Of course, ultralight dark matter could take on any mass a priori, and ground-based GW detectors only allow us to probe a small region of this parameter space, from $\sim 10^{-14}-10^{-12}$ eV. Space-based GW detectors will permit being sensitive to masses a few orders of magnitude smaller, while high-frequency gravitational-wave (HFGW) detectors could see DM particles a few orders of magnitude larger Aggarwal et al. (2021). However, it should be noted that when placing constraints on ultralight dark matter, the stochasticity of the signal could affect the limits: if we happen to “get unlucky” and observe the signal when the phase offset reduces the signal amplitude to zero, as can be seen at the zero points in Fig. 3, we would have no constraint. Refs. Nakatsuka et al. (2023); Boddy et al. (2025) have shown that for ground- and space-based interferometers, that this effect has an $\mathcal{O}(1)$ impact on upper limits, but needs to be mitigated for pulsar timing arrays by averaging over the signals coming from many pulsars.
@@
The prospects for future ground-based GW interferometers, such as ET and Cosmic Explorer, to observe ultralight dark matter interactions are bright. Within a decade or so of observation, such detectors could actually be competitive at masses between $10^{-15},10^{-14}]$ eV, thus permitting an extremely strong probe of new physics Pierce et al. (2018); Morisaki et al. (2021).
@@
Though GW interferometers can probe extremely small DM/standard model couplings, we do not have an estimate of a minimum strength of these interactions. We thus have no guarantee, in any search that we do, that we are getting closer to probing the true coupling value between ordinary matter and the dark sector. Such a caveat is balanced by the relatively model agnostic nature of our searches, which make minimal assumptions about the signal model, and essentially just look for correlated noise. While we probably cannot believe any one model for DM/standard model interactions, we can be sure that our techniques are sensitive to a wide range of theories, even ones that have yet to be thought of.
IV Macroscopic DM transiting through GW interferometers
As mentioned in Section I, the mass of DM could span many orders of magnitude. In particular, macroscopic dark matter, with masses in the range $M_{\rm DM}\sim[1,10^{9}]$ kg, has been proposed as a viable candidate Hall et al. (2018). These candidates could include composite objects such as nuclear-density nuggets Itoh (1970); Witten (1984) or Q-balls Coleman (1985),
If such objects were to transit through or near a GW interferometer, they could exert measurable gravitational forces on the test masses. In most models, the dominant interaction would be gravitational: the passage of a massive, compact object would cause a time-varying, localized gravitational acceleration of the test masses, leading to a transient displacement signal in the detector.
The resulting signature would be a short-duration, impulsive signal, potentially localized in time and space depending on the trajectory of the DM object. Unlike continuous-wave or stochastic DM signals discussed in Section II, this effect would appear as a transient, with a characteristic waveform determined by the mass, size, velocity, and trajectory of the object.
The event rate of such transits is inversely proportional to the mass of the dark matter particles, and the number density $n_{\rm DM}$ depends on $\rho_{\text{DM}}$:
$\rho_{\text{DM}}=M_{\rm DM}\cdot n_{\text{DM}},$ (79)
This equation implies that the characteristic length scale $L_{\rm DM}$ associated with such objects (assuming constant density), i.e. the distance between DM objects, is given by:
$\frac{L_{\rm DM}}{10^{4}~\text{km}}\simeq 1.2\times\left(\frac{M_{\text{DM}}}{1~\text{kg}}\right)^{1/3}.$ (80)
For $M_{\rm DM}=1$ kg, this corresponds to a flux of $\Phi\sim n_{\rm DM}v_{0}\sim 3\times 10^{-10}$ km-2 s-1, assuming a typical DM velocity $v_{0}\sim 220$ km/s. Moreover, the size of DM, $r_{\rm DM}$, would be tiny with respect to $L_{\rm DM}$: $r_{\rm DM}\sim(M_{\rm DM}/(\frac{4}{3}\pi\rho_{\rm nugget}))^{1/3}\sim 3\mu\text{m}$ for $\rho_{\rm nugget}\sim 10^{17}\text{ kg/m}^{3}$. This implies that one transient event per year would be expected in a detector with an effective cross-sectional area777the area in the plane perpendicular to the DM flux within which a passing object would produce a transient of order a few hundred km2, i.e. DM would have an impact parameter of $\sim 10$ km, which is of the same order as the length of the interferometer arms (4 km). However, gravitationally interacting DM would not be able to produce such an event rate Hall et al. (2018); Du et al. (2023), so we will need to consider interactions of DM with the standard model to enhance the event rate. In Section IV.1, we will discuss the physics of Yukawa-like interactions that macroscopic DM can have with the standard model, which will enhance the strength of the transiting DM object. Then, in Section IV.2, we show the projected constraints on Yukawa-like interactions using future data from ground-based GW interferometers.
IV.1 Yukawa-like interactions